OpenStudy (anonymous):

Find exact value if: 0

OpenStudy (anonymous):

Now I KNOW I answered this question yesterday AND a few minutes ago.

OpenStudy (anonymous):

what a answer?

OpenStudy (anonymous):

Do you know the formula for sin(a+b)?

OpenStudy (anonymous):

yes, but I not sure I solve it right

OpenStudy (anonymous):

Post how you solved it. I'll tell you if it's right, or where you went wrong.

OpenStudy (anonymous):

can you show your work?

OpenStudy (anonymous):

Can YOU show YOUR work?

OpenStudy (anonymous):

\[\sin\frac{4}{9}\cos\frac{1}{4}+\cos\frac{4}{9}\sin\frac{1}{4}\]

OpenStudy (anonymous):

x doesn't equal 4/9. sin x = 4/9. Use the sin(a+b) formula to expand sin(x-y)

OpenStudy (anonymous):

I know you know how to do it because you did it before.

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

What's the formula for sin(a+b)?

OpenStudy (anonymous):

\[cosx=\sqrt{1-(\frac{4}{9})^2}\]this I folow from the book

OpenStudy (anonymous):

That's right, but first you have to expand sin(x-y). To do that you need to know the rule for expanding sin(a+b). What is the rule?

OpenStudy (anonymous):

sin(a)cos(b)+cos(a)sin(b)

OpenStudy (anonymous):

Correct. Now apply that formula to sin(x-y). You use the rule you just typed, but set a=x and b=-y.

OpenStudy (anonymous):

sin(4/9)cos(-1/4) +cos(4/9)sin(-1/4)?

OpenStudy (anonymous):

Whoa! Hold on there. Don't start substituting values just yet. Expand sin(x-y) using the sin(a+b) formula you just typed.

OpenStudy (anonymous):

Also, x and y aren't 4/9 and 1/4. SIN x = 4/9 and SIN y = 1/4. There's a big difference.

OpenStudy (anonymous):

how you solve?

OpenStudy (anonymous):

Do what I asked you to do. Expand sin(x-y) using the formula you already posted for expanding sin(a+b). Use a=x and b=-y.

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

You wrote sin(a+b) = sin(a)cos(b) + cos(a)sin(b). Use that formula to expand sin(x-y). Just plug in x for a and -y for b.

OpenStudy (anonymous):

Post what you got.

OpenStudy (anonymous):

I did post

OpenStudy (anonymous):

sin(a+b) = sin(a)cos(b) + cos(a)sin(b). sin(4/9)cos(-1/4) +cos(4/9)sin(-1/4)?

OpenStudy (anonymous):

Copy and post it again, because I don't see it. It should be an expression in x and y.

OpenStudy (anonymous):

sin(a+b) = sin(a)cos(b) + cos(a)sin(b). sin(4/9)cos(-1/4) +cos(4/9)sin(-1/4)?

OpenStudy (anonymous):

Neither of those are an expression in x and y.

OpenStudy (anonymous):

Also, you don't know the values of x or y. You only know the values of sin(x) and sin(y).

OpenStudy (anonymous):

So you need to find an expression involving sin(x), sin(y), cos(x), and cos(y) so you can plug in the values that you DO know.

OpenStudy (anonymous):

That's why I keep asking you to use the sin(a+b) formula to expand sin(x-y). That will yield an expression involving sin(x), sin(y), cos(x), and cos(y).

OpenStudy (anonymous):

It getting late I need go sleep,I will work on it, I f I don't get I will post again tomorrow,ty

OpenStudy (anonymous):

sin(x - y) =sinx*cosy - cosx*siny

OpenStudy (anonymous):

\[sin^2\theta + cos^2\theta = 1 \]

OpenStudy (anonymous):

You can get cos values with the above identity